OpenStudy (anonymous):

Solve for x: −5|x + 1| = 10

OpenStudy (nurali):

−5|x + 1| = 10 |x + 1|=10/-5 |x + 1|=-2 x+1=-2 or x+1=-(-2) x=-3 or x=1

OpenStudy (nurali):

Welcome to Openstudy.

OpenStudy (anonymous):

thank you so much!! @Nurali

OpenStudy (nurali):

Anytime.

OpenStudy (nurali):

Medal Please.